mitch1738:

Solve for y. d = 5x + 5y y=d−5xx y=d−5x5 y=5(d−5x) y=5x−d5

dude:

okay so first get x to the other side, do you know how to do that?

mitch1738:

im special ed i dont knowhow to do it

dude:

Okay what is the opposite of addition?

mitch1738:

subtract

mitch1738:

so can you just give me the answer plz

mitch1738:

i have an f

dude:

Site does not allow me to give direct answers but I can assist you in getting the answer \(~~~~~d = 5x + 5y\\ -5x~~-5x\) Subtract 5x on both sides

mitch1738:

thats zero bruh lol

dude:

On the right side yes \(-5x+d=5y\)

mitch1738:

so this the answer

dude:

Not quite now we have to isolate 5 from the y

dude:

What is the opposite of multiplication?

mitch1738:

divide

mitch1738:

just give me the damn answer please

mitch1738:

im sorry i snapped

mitch1738:

its just so hard i need the answer now

dude:

So divide 5 on both sides \(−5x+d=5y\\ \overline{ ~~~~~~ 5~~~~~}~~~~~~ \overline{5}\)

dude:

\(−5x+d=y\\ \overline{ ~~~~~~ 5~~~~~}\)

dude:

wow I really screwed there my bad

dude:

Anyway that is how you got to the answer